OpenStudy (anonymous):

what is the systems equation of 2x + 6y = 4

OpenStudy (solomonzelman):

I only see one equation.

OpenStudy (agreene):

A) \[2x+6y=4\] \[\frac{2x+6y=4}{2}\] \[x+3y=2\] \[x=2-3y\] B) \[2x+6y=4\] \[\frac{2x+6y=4}{6}\] \[\frac{x}{3}+y=\frac23\] \[y=\frac23-\frac{x}3\] So the sytem would be \[x=2-3y\] \[y=\frac23-\frac{x}3\]